# Robustness of deep learning classification to adversarial input on GPUs: asynchronous parallel accumulation is a source of vulnerability

### v1

@@ 2.3.2 Learnable Permutation to Find Possible Adversarial Perturbations
We propose a gradient-based optimization technique to find a permutation of floating point operations that causes misclassification. Following Section 2.1, let $f$ be a classifier mapping an input tensor $\mathbf{x}$ to logits, a probability vector of length $L$, the number of classes. We take the argmax of the logits to obtain the final classification. We require $f$ to include floating point accumulations. Due to APFPR, the output of $f$ depends on a permutation matrix $P$ describing the order reductions, written as $f(P,\mathbf{x})$. In cases where $f$ is composed of multiple functions with the same properties, we parameterize using a set of permutation matrices $\{P_{i}\}$. The classifier is trained by minimizing a loss function $\mathcal{L}(f,y)$, where $y$ is the ground truth label. To find the $\{P_{i}\}$ which causes $f$ to misclassify $\mathbf{x}$, we maximize the prediction error with respect to the permutation perturbation:
@@
We use the Gumbel-Softmax technique [12, 5] to create a differentiable approximation of the permutations matrices. By adding Gumbel noise and applying softmax to the matrix, we can use gradient descent to optimize the set $\{P_{i}\}$ that maximizes the loss function, with the other parameters of $f$ fixed. Next, valid permutation matrices are obtained by solving the linear assignment problem via the Hungarian algorithm [9]. This method, inspired by adversarial attacks (Section 2.1), maximizes error with respect to floating-point operation ordering instead of the input. While the approach does not guarantee misclassification, it provides a more efficient way to find adversarial permutations compared to brute force search.
We now present practical steps to use the LP method: (1) Identify non-deterministic functions by referencing documentation or using a linter like the torchdet tool [13], (2) For each non-deterministic function, introduce a permutation matrix $P$ to simulate runtime variations in reductions. For example, consider a fully connected linear layer with a weight matrix $w$ of size $N\times M$ and a bias vector $\mathbf{b}$ of size $N$, where the intermediate output $\mathbf{y}$ is given by $\mathbf{y}=w^{T}\mathbf{x}+\mathbf{b}$. To simulate variation in accumulation order, we use the permutation matrix $P$ to permute element-wise products, $S_{i}$ before reduction where $S_{i}=\{w_{i0}x_{0},\dots,w_{iM}x_{M}\}$, computing $y_{i}=\sum_{j=1}^{M}(P\cdot S_{i})_{j}$. For the linear classifier in Sec 2.2.1, we introduce the permutation matrix $P$ as follows: $P\hat{\mathbf{n}}\cdot P\mathbf{x}$. (3) Perform gradient descent as specified in Eq. 5, optimizing only over permutation matrices. (4) Perform a forward pass and mark any misclassifications to generate a worst-case bound. As shown in the right panel of Fig 1, the LP approach provides a tight bound to the EWA attack. Next, we investigate misclassifications in GNNs, extending prior work [19] which identified significant run-to-run output variability in GNNs but did not consider misclassification.
@@ 3 Non-determinism in Graph Neural Networks
Graph neural networks (GNNs) operate on unordered graph data. For a graph $G=(V,E)$, any permutation of $V$ and $E$ represents the same structure. GNNs learn node and edge representations via message passing and aggregation, the core operations in most architectures [27]. Since node neighborhoods lack a fixed order, GNNs rely on permutation-invariant aggregation like add and mean implemented in PyTorch Geometric [1] with scatter_reduce functions, which introduce non-determinism due to atomic operations. This, combined with the non-unique representation of graphs, makes GNNs in PyTorch Geometric well-suited for studying APFPR effects.
We investigate run-to-run variability in robustness results, identifying worst-case accuracies with the learnable permutation approach. Additionally, we perform the EWA attack (Section 2.3.1) to induce misclassifications and evaluate the ability of the LP approach to provide worst-case estimates.
@@ 3.1 Experimental Methodology
We study APFPR vulnerability in the GNNS architectures: GraphSAGE, GAT, and GCN [4, 24, 8], using the CORA, CiteSeer, and PubMed datasets [16]. These widely used benchmarks evaluate GNN performance in semi-supervised node classification. For each model-dataset pair, we analyze misclassifications due to non-deterministic functions and EWA. We train $N_{\text{train}}=100$ models for 25 epochs, initializing them identically with fixed randomness from stochastic training and random seed settings. Training models with atomic functions have been shown to produce different weights due to APFPR [19] and we aim to investigate the full training and inference pipeline.
To assess inference variability, we perform $N_{\text{val}}=10,000$ forward passes on the validation set, with and without atomics. Note the base class of PyTorch GNNs are non-deterministic by default [19]. To provide a deterministic control experiment, we refactor the base class with a deterministic index_add operation replacing scatter_reduce. However this is not a solution to the non-determinism problem in GNNs since significant refactoring is needed to ensure both functional parity and sufficient performance and it is unclear if this is possible. An input is marked misclassified if any iteration produces an incorrect prediction. A large $N_{\text{val}}$ is required since misclassifications may only appear after many identical repeated runs, as shown in Section 2.2.1. We prevent kernel switching, isolating APFPR as the sole source of classification flips. We also predict misclassifications and worst-case accuracy bounds using the LP method (Section 2.3.2) with $1,000$ optimization steps. Graph edges in PyTorch Geometric are permutation-invariant, allowing us to introduce floating-point accumulation sensitivity through permutation matrices in GNN layers when passing the adjacency matrix or edge_index variable across layers.
Table 1: Average accuracy (number of correct classifications out of 500) on the CORA dataset for a 10-layer GraphSAGE model under different attacks and attack epsilon values, with standard deviation. "ND" and "D" indicate non-deterministic or deterministic PyTorch settings during inference, respectively. For $ND$, 10,000 inference runs are performed. "LP" refers to a learnable permutation worst-case bound, determined for each input and "EWA" refers to the external workload attack, which succeeds at least 75% of the time on 1,000 repeated runs. We bold experiments which result in misclassifications. All experiments are performed on the H100 with default PyTorch FP32 precision.
@@
Additionally, we perform the EWA to examine whether external workloads can induce misclassification in a real-world network. As before, we run the Bayesian optimization method for 100 iterations, conducting 1,000 experiments per iteration, to determine the optimal attack matrix size $k\in\{1,000,10,000\}$ for flipping classifications to the second most probable class. We then perform 1,000 repeated inferences to assess attack success, considering it successful if misclassification occurs at least 75% of the time. While this threshold is arbitrary, a reliable attack should consistently induce misclassification.
We validate models and test both unperturbed and adversarial inputs from FGSM, PGD, random, and targeted attacks (Section 2.1).
@@ 3.2 Results
The results of experiments for a 10-layer GraphSAGE model on the CORA dataset are shown in Table 1, performed on a H100. Similar behavior was observed across other datasets, models and GPUs (details available on our GitHub). For each adversarial attack method and epsilon value, we report test accuracy as the number of correct classifications (out of 500), with the first row showing results for $\epsilon=0$ (no attack). Columns labeled $ND$ or $D$ indicate whether deterministic PyTorch kernels were used during inference, and the LP column represents worst-case upper bounds determined via learned permutation optimization. The EWA column represents the EWA attack, which use naive matrix multiplication workloads. Errors are standard deviations over the $100$ trained models.
As expected, accuracies decrease with increasing attack strength $\epsilon$. Toggling PyTorch’s non-deterministic functions has little impact on average accuracy (D and ND columns). However, adversarial accuracy varies significantly at certain epsilon values, indicating that APFPR induces additional misclassifications beyond input perturbations. Notably, large errors occur even at $\epsilon=0$, showing that non-perturbed inputs are vulnerable to APFPR. EWA reliably misclassifies such inputs, with targeted and PGD attacks being the most affected—leading to adversarial accuracy drops of up to 4.6% (Targeted Attack, $\epsilon=1e-2$). Random attacks are minimally impacted. The EWA attack works $\geq$ 75% of the time, which is at least a three order magnitude increase in run-to-run misclassification consistency. EWA achieves convergence in under 80 iterations during the optimization step. Tightening the constraint $\geq$ 85% makes the EWA ineffective and relaxing it $\leq$ 20% makes it always effective, producing accuracy values closer to the LP approach. The LP method provides strong worst-case bounds and should be integrated into existing robustness verification tools. A workload as simple and common as a matrix multiplication in the EWA is sufficient to induce misclassifications in GNNs and simpler linear classifiers as in Sec. 2.3.1. As before, an erratic trend between optimal matrix size and inputs was found. We leave an in depth analysis to future work, requiring the development of novel GPU scheduler probing tools.
@@ 4 Impact of GPU State on Order of Reductions
The previous section demonstrates that GPU states can significantly impact ML workloads and should be considered as seriously as more conventional attacks. However, it only provides an indirect measure of the mechanisms behind misclassifications. In this section, we directly measure how the GPU state affects the ordering of asynchronous parallel operations, such as those involving CUDA’s atomicAdd. While the exact scheduling of operations on the GPU is not yet fully understood [15, 14] , it plays a critical role in determining the order of these operations. To our knowledge, no prior studies have specifically explored how the GPU state influences the ordering of asynchronous operations.
@@ 4.1 Methodology
We use the parallel sum algorithm to show how the asynchronous operation order, measured by block index vs execution order (BIEO), is influenced by external loads. The reduction $\sum_{i}^{n}x_{i}$, where $x_{i}\neq x_{j}\ \forall i,j,\ x_{i}>0$ double precision floating point numbers (FP64) numbers, runs on a GPU using atomicAdd, which has an undefined execution order. To recover the execution order, we track the accumulator updates per block and sort them post-execution. We sum 20 lists of one million uniform FP64 numbers, with and without an additional double-precision matrix-matrix multiplication (DGEMM) workload. Each test runs 10 times to account for system variations. Sorting yields two datasets—reduction only (RO) and with a DGEMM running in a different cuda stream (RDGEMM). The Kendall’s $\tau$ correlation[7] measures permutation similarity: $K_{\textsc{RO}}$ for RO and $K_{\textsc{RO-RDGEM}}$ comparing RO to RDGEMM. Further work probing the GPU scheduler is necessary to investigate reductions of more general arrays.
Figure 2: Left panel: PDF $K_{\textsc{RO}}$ and $K_{\textsc{RO-RDGEM}}$ on GH200. Right panel: $K_{\textsc{RO-RDGEM}}$ PDF for six MiG configurations on GH200. The inset in the left figure shows BIEO for the RO and RDGEMM workloads with the lowest $K_{\textsc{RO-RDGEM}}$ correlation.
@@ 4.2 Results
This method shows the GPU states effects on the execution order, aiding with verification in non-deterministic settings. We test various GPUs (Section 6), power settings, and partitions, presenting results for GH200 and V100.
@@ 4.2.1 Impact of external workloads
We calculated $K_{\textsc{RO}}$ and $K_{\textsc{RO-RDGEM}}$ for the GH200 GPU architecture. As shown in Fig. 2, GH200 exhibits distinct block scheduling and atomic instruction behaviors. The $K_{\textsc{RO}}$ distribution ranges from 0.31 to $0.70$, peaking at 0.66, while $K_{\textsc{RO-RDGEM}}$ ranges from $0.32$ to $0.73$ with a mean of $0.67$. The bimodal distributions for DGEMM workloads differ from the uni-modal distributions for unperturbed reductions, reflecting sensitivity to external workloads. As shown on Fig. 2, the V100 GPU behave differently as $K_{\textsc{RO}}$ and $K_{\textsc{RO-RDGEM}}$ distributions are narrower than on GH200. The $K_{\textsc{RO-RDGEM}}$distribution remains mult-modal but has a more complex structure than on the GH200 GPU. The inset highlights the BIEO snapshot for the pair with the lowest Kendall $\tau$ correlation, showing non-sequential block index ordering at first, which later converges into two parallel distributions. These results show that external workloads significantly influence the execution order of atomicAdd on GH200, expanding the range of possible ordering permutations.
Figure 3: Left panel: $K_{\textsc{RO-RDGEM}}$ for six different MiG configurations on GH200 showing the effect of resource restrictions on the Kendall $\tau$ correlations. Right panel: impact of power capping on the Kendall $\tau$ correlations.
@@ 4.2.2 Impact of MiG Configuration
The GH200 supports up to seven MiG partitions, enabling efficient resource sharing. The left panel of Fig. 3 shows $K_{\textsc{RO-RDGEM}}$ across MiG configurations. Higher values occur with fewer SM units and decrease as resources increase while wider distributions in larger GPUs suggest greater permutation variability. Since MiG partitions share the same hardware, each slice’s behavior depends on the runtime configuration. This is especially relevant in virtualized cloud environments.
@@ 4.2.3 Impact of Power
Research shows that power capping improves energy efficiency and reduces hardware failure rates in HPC and DL workloads [26]. Modern GPUs, like the NVIDIA GH200, include power capping to regulate power draw [26]. While power capping reduces GPU clock speeds, increasing scheduling latency and slowing thread execution, it can also lead to resource contention and impact the atomic operation ordering in compute-intensive tasks. As shown on the right panel of Fig.3 we observed no significant differences in the PDF of $K_{\textsc{RO}}$ and $K_{\textsc{RO-RDGEM}}$, suggesting that power capping does not notably affect instruction ordering.

### v2

@@ 2.3.2 Learnable Permutation to Find Possible Adversarial Perturbations:
We developed a gradient-based optimization technique to find a permutation of floating-point operations that causes misclassification. Following Section 2.1, let $f$ be a classifier mapping an input tensor $\mathbf{x}$ to logits a probability vector of length $L$, the number of classes; in some cases $f$ may be composed of multiple functions $f_{i}$ with the same properties. We take the argmax of the logits to obtain the final classification. We require $f$ to include floating-point accumulations. Due to APFPR, the output of $f$ depends on a set of permutation matrices $P_{i}$ describing the order of reductions, written as $f(P_{i},\mathbf{x})$, where $i$ is the index for each permutation matrix associated to the functions $f_{i}$ composing the function $f$. The classifier is trained by minimizing a loss function $\mathcal{L}(f,y)$, where $y$ is the ground truth label. To find $\{P_{i}\}$ that cause $f$ to misclassify $\mathbf{x}$, we maximize the prediction error with respect to the permutation perturbation:
@@
We use the Gumbel-Softmax technique [12, 5] to create a differentiable approximation of the permutation matrices. By adding Gumbel noise and applying softmax to the matrix, we can use gradient descent to optimize the set $\{P_{i}\}$ that maximizes the loss function, with the other parameters of $f$ fixed. Next, valid permutation matrices are obtained by solving the linear assignment problem via the Hungarian algorithm [9]. This method, inspired by adversarial attacks (Section 2.1), maximizes error with respect to floating-point operation ordering instead of the input. While the approach does not guarantee misclassification, it provides a more efficient way to find adversarial permutations compared to brute-force search.
We now present practical steps to use the LP method: (1) Identify non-deterministic functions by referencing documentation or using a linter like the torchdet tool [13], (2) For each non-deterministic function, introduce a permutation matrix $P$ to simulate runtime variations in reductions. For example, consider the linear classifier in Sec. 2.2.1. In that case, we compute $P\hat{\mathbf{n}}\cdot P\mathbf{x}$ instead of $\hat{\mathbf{n}}\cdot\mathbf{x}$. For a fully connected linear layer with a weight matrix $w$ of size $N\times M$ and a bias vector $\mathbf{b}$ of size $N$, where the intermediate output $\mathbf{y}$ is given by $\mathbf{y}=w^{T}\mathbf{x}+\mathbf{b}$, we apply the permutation matrix $P$ on element-wise products $S_{i}$ before reduction where $S_{i}=\{w_{i0}x_{0},\dots,w_{iM}x_{M}\}$, computing $y_{i}=\sum_{j=1}^{M}(P\cdot S_{i})_{j}$. (3) Perform a gradient descent as specified in Eq. 5, optimizing only over permutation matrices. (4) Perform a forward pass and mark any misclassifications to generate a worst-case bound. As shown in the right panel of Fig. 1, the LP approach provides a tight bound on the EWA attack. Next, we investigate misclassifications in GNNs, extending previous work [19] which identified significant run-to-run output variability in GNNs but did not consider misclassification.
@@ 3 Non-Determinism in Graph Neural Networks
Graph neural networks (GNNs) operate on unordered graph data. For a graph $G=(V,E)$, any permutation of $V$ and $E$ represents the same structure. GNNs learn node and edge representations via message passing and aggregation, the core operations in most architectures [27]. Since node neighborhoods lack a fixed order, GNNs rely on permutation-invariant aggregation like add and mean implemented in PyTorch Geometric [1] with scatter_reduce functions, which introduce non-determinism due to atomic operations. This, combined with the non-unique representation of graphs, makes GNNs in PyTorch Geometric well-suited for studying APFPR effects. On these GNNs we investigate run-to-run variability in robustness results, identifying worst-case accuracies with the learnable permutation approach. Additionally, we perform the EWA attack (Section 2.3.1) to induce misclassifications and evaluate the ability of the LP approach to provide worst-case estimates.
@@ 3.1 Experimental Methodology
We study APFPR vulnerability in GNN architectures: GraphSAGE, GAT, and GCN [4, 24, 8], using the CORA, CiteSeer, and PubMed datasets [16]. These widely used benchmarks evaluate GNN performance in semi-supervised node classification. For each model-dataset pair, we analyze misclassifications due to non-deterministic functions and EWA. We train $N_{\text{train}}=100$ models for 25 epochs, initializing them identically with fixed randomness from stochastic training and random seed settings. Training models with atomic functions have been shown to produce different weights due to APFPR [19] and we aim to investigate the full training and inference pipeline. To assess inference variability, we perform $N_{\text{val}}=10000$ forward passes on the validation set, with and without atomics. The base class of PyTorch GNNs is non-deterministic by default [19]. To provide a deterministic control experiment, we refactor the base class with a deterministic index_add operation replacing scatter_reduce. However, this is not a solution to the non-determinism problem in GNNs since significant refactoring is needed to ensure both functional parity and sufficient performance and it is unclear if this is possible. An input is marked as misclassified if any iteration produces an incorrect prediction. A large $N_{\text{val}}$ is required since misclassifications may only appear after many identical repeated runs, as shown in Section 2.2.1. We prevent kernel switching, isolating APFPR as the sole source of classification flips. We also predict misclassifications and worst-case accuracy bounds using the LP method (Section 2.3.2) with $1000$ optimization steps. Graph edges in PyTorch Geometric are permutation invariant, allowing us to introduce floating-point accumulation sensitivity through permutation matrices in GNN layers when passing the adjacency matrix or edge_index variable across layers.
Table 1: Average accuracy (number of correct classifications out of 500) on the CORA dataset for a 10-layer GraphSAGE model under different attacks and attack epsilon values, with standard deviation. “ND” and “D” indicate non-deterministic or deterministic PyTorch settings during inference, respectively. For NDND, 10000 inference runs are performed. “LP” refers to a learnable permutation worst-case bound, determined for each input and “EW” refers to the external workload attack, which succeeds at least 75% of the time on 1000 repeated runs. We bold experiments which result in misclassifications. All experiments are performed on the H100 with default PyTorch FP32 precision.
@@
Additionally, we performed EWA to examine whether external workloads can induce misclassification in a real-world network. As before, we ran Bayesian optimization for 100 iterations (with 1000 experiments per iteration) to identify the optimal attack matrix size $k\in\{1000,10000\}$ that flips classifications to the second most probable class. We then perform 1000 repeated inferences to assess the attack’s success, considering it successful if misclassification occurs at least 75% of the time. While this threshold is arbitrary, a reliable attack should consistently induce misclassification. We validate models and test unperturbed and adversarial inputs from FGSM, PGD, random, and targeted attacks (Section 2.1).
@@ 3.2 Results
Results for a 10-layer GraphSAGE model on the CORA dataset are shown in Table 1, performed on an H100. Similar behavior was observed in other datasets, models, and GPUs (details are available on our GitHub). For each adversarial attack method and epsilon value, we report test accuracy as the number of correct classifications (out of 500), with the first row showing results for $\epsilon=0$ (no attack). The Columns labeled $ND$ or $D$ indicate whether deterministic PyTorch kernels were used during inference, and the LP column represents worst-case upper bounds determined via learned permutation optimization. The EWA column represents the EWA attack, which use naive matrix multiplication workloads. Errors are standard deviations over the $100$ trained models.
As expected, accuracies decrease with increasing attack strength $\epsilon$. Toggling PyTorch’s non-deterministic functions on or off has little impact on average accuracy (D and ND columns); not all PyTorch functions have a deterministic version [19]. However, adversarial accuracy varies significantly at certain epsilon values, indicating that APFPR induces additional misclassifications beyond input perturbations. Notably, large errors occur even at $\epsilon=0$, showing that non-perturbed inputs are vulnerable to APFPR. EWA reliably misclassifies such inputs, with targeted and PGD attacks being the most affected, leading to adversarial accuracy drops of up to 4.6% (Targeted Attack, $\epsilon=0.01$). Random attacks are minimally impacted. The EWA attack works $\geq$ 75% of the time, which is at least a three-order-of-magnitude increase in run-to-run misclassification consistency. EWA achieves convergence in less than 80 iterations during the optimization step. Tightening the constraint $\geq$ 85% makes EWA ineffective and relaxing it $\leq$ 20% makes it always effective, producing accuracy values closer to the LP approach. The LP method provides strong worst-case bounds and should be integrated into existing robustness verification tools as a workload as simple and common as a matrix multiplication in EWA is sufficient to induce misclassifications in GNNs and simpler linear classifiers as in Sec. 2.3.1. As before, an erratic trend was found between optimal matrix size and inputs. We leave an in-depth analysis to future work, requiring the development of novel GPU scheduler probing tools.
@@ 4 Impact of GPU State on Order of Reductions
The previous section demonstrates that GPU states can significantly impact ML workloads and should be considered as seriously as more conventional attacks. However, it only provides an indirect measure of the mechanisms behind misclassifications. In this section, we directly measure how the GPU state affects the ordering of asynchronous parallel operations, such as those involving CUDA’s atomicAdd. While exact scheduler behavior is not always understood [15, 14], the scheduler plays a critical role in determining the order of these operations. To our knowledge, no prior studies have specifically explored how the GPU state influences the ordering of asynchronous operations.
@@ 4.1 Methodology
We use the parallel sum algorithm to show how the asynchronous operation order, measured by block index vs execution order (BIEO), is influenced by external loads. The reduction $\sum_{i}^{n}x_{i}$, where $x_{i}\neq x_{j}\ \forall i,j,\ x_{i}>0$ double precision floating-point numbers (FP64) numbers, runs on a GPU using atomicAdd, which has an undefined execution order. To recover the execution order, we track the accumulator updates per block and sort them post-execution. We sum 100 lists of 1M uniform FP64 numbers, with and without an additional double-precision matrix-matrix multiplication (DGEMM) workload. Each test was run 10 times to account for system variations. Sorting yields two datasets—reduction only (RO) and with a DGEMM running in a different CUDA stream (RDGEMM). The Kendall’s $\tau$ correlation [7] measures permutation similarity: $K_{\textsc{RO}}\$ for RO and $K_{\textsc{RO-RDGEM}}$ comparing RO to RDGEMM. This method shows the GPU states’ effects on the execution order, aiding with verification in non-deterministic settings. We tested various GPUs (Section 6), power settings, and partitions, presenting results for GH200 and V100.
@@ 4.2.1 Impact of External Workloads
We calculated $K_{\textsc{RO}}\$ and $K_{\textsc{RO-RDGEM}}$ for the GH200 GPU architecture. As shown in Fig. 2, GH200 exhibits distinct block scheduling and atomic instruction behaviors. The $K_{\textsc{RO}}\$ distribution ranges from $0.32$ to $0.70$, peaking at $0.67$, while $K_{\textsc{RO-RDGEM}}$ ranges from $0.45$ to $0.83$ with a mean of $0.71$. The multimodal distributions for DGEMM workloads differ significantly from the unimodal distributions for unperturbed reductions, reflecting the sensitivity to external workloads.
Figure 2: Left panel: PDF $K_{\textsc{RO}}\$ and $K_{\textsc{RO-RDGEM}}$ on GH200. Right panel: $K_{\textsc{RO-RDGEM}}$ PDF $K_{\textsc{RO}}\$ and $K_{\textsc{RO-RDGEM}}$ on V00. The inset in the figures shows BIEO for the RO and RDGEMM workloads with the lowest $K_{\textsc{RO-RDGEM}}$ correlation.
As shown in Fig. 2, the V100 GPU behaves differently, as $K_{\textsc{RO}}\$ and $K_{\textsc{RO-RDGEM}}$ distributions are narrower than on GH200. The $K_{\textsc{RO-RDGEM}}$ distribution remains multimodal but has a more complex structure than on the GH200 GPU. The inset highlights the BIEO snapshot for the pair with the lowest Kendall $\tau$ correlation, showing non-sequential block index ordering at first, which later converges into two parallel distributions. These results show that external workloads significantly influence the execution order of atomicAdd on GH200, expanding the range of possible ordering permutations.
@@ 4.2.2 Impact of MiG Configuration
The GH200 supports up to seven multi-instance GPU (MiG) partitions, allowing resource sharing. The left panel of Fig. 3 shows $K_{\textsc{RO-RDGEM}}$ across MiG configurations. Higher values occur with fewer SM units and decrease as resources increase, while wider distributions in larger GPUs suggest greater permutation variability. Since MiG partitions share the same hardware, each slice’s behavior depends on the runtime configuration. This is especially relevant in virtualized cloud environments.
Figure 3: Left panel: $K_{\textsc{RO-RDGEM}}$PDF for six different MiG configurations on GH200 showing the effect of resource restrictions on the Kendall $\tau$ correlations. Right panel: impact of power capping on the Kendall $\tau$ correlations.
@@ 4.2.3 Impact of Power
Research shows that power capping improves energy efficiency and reduces hardware failure rates in HPC and DL workloads [26]. Modern GPUs, such as the NVIDIA GH200, include power capping to regulate power draw [26]. While power capping reduces GPU clock speeds, increasing scheduling latency and slowing thread execution, it can also lead to resource contention and impact the atomic operation ordering in compute-intensive tasks. As shown in the right panel of Fig.3 we observed no significant differences in the PDF of $K_{\textsc{RO}}\$ and $K_{\textsc{RO-RDGEM}}$, suggesting that power capping does not notably affect instruction ordering.